# Condense to Conduct and Conduct to Condense

### v1

@@ Abstract
In this paper we give the first examples of low-conductance permutations. The notion of conductance of permutations was introduced in the paper ”Indifferentiability of Confusion-Diffusion Networks” by Dodis et al., where the search for low-conductance permutations was initiated and motivated. In this paper we not only give the desired examples, but also make a general characterization of the problem – i.e. we show that low-conductance permutations are equivalent to permutations that have the information-theoretic properties of the so-called Multi-Source-Somewhere-Condensers.
@@ 1.1 Background
The starting concept of this paper is the notion of permutation conductance. This notion was introduced by Yevgeniy Dodis, Martijn Stam, John Steinberger, and Tianren Liu in their seminal paper Indifferentiability of confusion-diffusion networks [11].
The above-mentioned paper concerns the theoretical analysis of the indifferentiability security of the confusion-diffusion networks, investigating how many rounds of certain constructions of block ciphers and hash functions provide adequate security. It turns out that in these considerations, the key primitive is a certain permutation, from which we expect a possibly small so-called conductance. Much more on the subject of motivation and direct relations between permutations with an appropriate conductance and the security of the studied ciphers based on confussion-diffusion networks can be found in [11]. In our work, the conductance of permutation is studied primarily as a combinatorial concept.
@@
From the formal definition side, each permutation of the type $\pi:\{0,1\}^{nw}\to\{0,1\}^{nw}$ has a certain degree of conductance that additionally depends on a certain parameter $\alpha\in(0,1)$. This degree (denoted $\textrm{CondDeg}_{\alpha}(\pi)$) is always a real number between $1$ and $w$ and is a kind of measure of how random $\pi$ is (the more random the permutation, the closer $\textrm{CondDeg}_{\alpha}$ is to $1$; in particular, a constant permutation has $\textrm{CondDeg}_{\alpha}$ equal to $w$). It turns out that permutations with low $\textrm{CondDeg}_{\alpha}$ (”more random”) are useful in the applications described in [11]. Fortunately, the authors of [11] showed (Theorem 3, App. B, probabilistic method) that almost every random permutation has (for parameters $\alpha<\frac{1}{2}-\frac{1}{nw}$, which is sufficient for the applications described in the paper) $\textrm{CondDeg}_{\alpha}$ very close to $1$ (more precisely: it is no larger than $1+\frac{\log(3nw)}{\alpha n}$). On the other hand – there were no explicit constructions showing how such permutations could be constructed in practice.
@@
In this paper we show for the first time direct examples of permutations with conductance degree less than $w$. Details are given in Theorem 1. Our constructions use the so-called Multi-Source-Somewhere-Condenser permutations, i.e. a certain – defined in this paper as Definition 5 – variant of classical somewhere-condensers (such as defined for exampe in Definition 1.11 in [18]).
@@
Interestingly, we also managed to obtain the inverse result – i.e. we show that every permutation with a sufficiently low conductance degree must satisfy the Multi-Source-Somewhere-Condenser assumptions. In other words, in this paper we attempt a full characterization of the problem, which is successful for almost the entire range of parameters (see Section 3.2 for details). In particular: this means that unfortunately we cannot count on low-conductance permutations expressed as, for example, simple linear functions, as hypothesized in [11].
To conclude the introduction, it is worth noting how the considered problem shows the non-obvious interpenetration of various fields in theoretical computer science and mathematics. Namely, the original problem can be classified as an example of research in the field of theoretical analysis of symmetric ciphers (which is not common in itself, see e.g. [17, 16, 12, 8]). A nice combinatorial problem was extracted from it (and it is described as such in the original paper). Then it turned out that its analysis leaned towards information theory (in particular the theory of min-entropy extraction), so it can be seen as a work in this field (and thus among papers such as [5, 7, 20, 15, 14, 22, 19, 9, 10]). It is also worth emphasizing that other similar reasonings, i.e. solving combinatorial problems using theorems from extractor theory, are not new, this phenomenon can also be observed e.g. in [13, 21, 2, 3, 4, 1]. This all shows an insight into why this problem turned out to be surprisingly non-trivial.
@@ 1.2.1 Main Notion
Since the above notion is exceptionally essential for this work, to increase its readability we add an additional, intuitive understanding of this notion below preceded by one more improtant definition:
@@
In simple terms, $\textrm{CondDeg}_{\alpha}(\pi)$ determines how much a given permutation $\pi$ distorts all discrete $q$-boxes of dimension $w$ for $q=2^{\alpha n}$. More specifically, for each such $q$-box $U$ we look at its image $\pi(U)$ and look for another $q$-box $V$ such that the intersection $U\cap V$ is as large as possible. The maximum we can obtain in this way is given by $\textrm{CondDeg}_{\alpha}$ in the form of the following relation:
@@
In the ”worst” (least random) case we will find such $U$ that $\pi(U)$ is just a another $q$-box $V$ and then $|\pi(U)\cap V|=q^{w}$, which gives $\textrm{CondDeg}_{\alpha}=w$.
On the other hand, ”more random” permutations should intuitively behave in such a way that regardless of the choice of $U$ and $V$, the cardinality of the set $\pi(U)\cap V$ will always be much smaller than $q^{w}$. At the same time, it is easy to see that it is always possible to choose $V$ such that $|\pi(U)\cap V|$ is at least $q$. In particular, this means that we definitely have $\textrm{CondDeg}_{\alpha}(\pi)\geq 1$.
@@ 1.2.2 Main result
In this paper we give the first example of a permutation with nontrivial conductance degree:
@@
If $n>n_{0}$ is prime and $\epsilon\leq\alpha\leq 0.8$, then there exists a permutation $\pi:\{0,1\}^{wn}\to\{0,1\}^{wn}$ such that
$\textrm{CondDeg}_{\alpha}(\pi)\leq w-\left\lfloor\frac{w}{3}\right\rfloor c,$
@@
The permutation $\pi:\{0,1\}^{wn}\to\{0,1\}^{wn}$ will be called a ($\alpha,\epsilon_{1},\epsilon_{2}$)-Multi-Source-Somewhere-Condenser if for any independent random variables $X_{1},X_{2},\ldots X_{w}\in\{0,1\}^{n}$ with min-entropy at least $\alpha n$, the distribution $\pi(X_{1},\ldots,X_{w})$ can be expressed as a convex combination of the distributions ($\sum_{1\leq i\leq w}\gamma_{i}C_{i}+\gamma R)$, such that $\gamma\leq\epsilon_{2}$ and $C_{i}]_{i}$ (i.e. the $i$-th coordinate of $C_{i}$) has min-entropy at least $(1+\epsilon_{1})\alpha n$.
Note: the definition above is based on Definition 1.11 from [18]. However we consider several sources of the input and we limit ourselves to the study of permutations. Nevertheless, its intuitive meaning remains the same: we define a transformation, which from several sources with smaller min-entropy, leads to a convex combination of distributions, each of which (except for one component of size $\gamma$) has min-entropy on some coordinate by a factor $(1+\epsilon_{1})$ larger.
Armed with the above definitions, we are ready to formulate the key technical lemma, the proof of which can be found in Section 2.1:
@@
The formula in the above lemma means that if $\epsilon_{2}$ is negligable, then we will achieve $\textrm{CondDeg}_{\alpha}\approx w-\epsilon_{1}$.
So, to effectively use Lemma 2, we need a Multi-Source-Somewhere-Condenser with suitably strong parameters, which fortunately exists and may be found for example in the paper [6]. Specifically, it holds (in [6] this result and its proof are discribed as Lemma 3.14):
@@
If $\epsilon\leq\alpha\leq 0.8$ and $n$ is prime, then the permutation $\pi$ defined above is ($\alpha,\delta,2^{-\delta\alpha n}$)-Multi-Source-Somewhere-Condenser.
@@
If $\epsilon\leq\alpha\leq 0.8$, then there exists a permutation $\pi:\{0,1\}^{3n}\to\{0,1\}^{3n}$ such that
$\textrm{CondDeg}_{\alpha}(\pi)\leq 3-c\text{ (for some }c>0\text{),}$
@@ 1.2.3 Converse Theorem
The proof is in Section 2.2.
@@ 1.3 Intuition and Toy Examples
Consider the function $\pi_{1}(a,b,c)=(a,b,c+ab)$, where $a,b,c\in GF(2^{n})$ and $n$ is a prime number. This is of course a permutation, since for any value of $(x,y,z)$, its only inverse image is (x,y,zxy)(x,y,z-xy). We want to justify why such defined $\pi_{1}:\{0,1\}^{3n}\to\{0,1\}^{3n}$ has conductance degree $\textrm{CondDeg}_{\alpha}(\pi_{1})$ (for some fixed parameter $\alpha$) strictly less than $3$.
@@
To achieve our goal, we will use the fact that $\pi_{1}$ is a Multi-Source-Somewhere-Condenser (thus Lemma 3), and even a stronger fact (which is formulated in this way in [6]), i.e. that it is always the third (not some) coordinate of $\pi_{1}(a,b,c)$ that is the condenser, except for some negligable part of the domain. More specifically, thanks to this version of Lemma 3 we know that – since we are considering a uniform distribution on $U$ – and thus we have that $U_{1},U_{2}$ and $U_{3}$ are independent and have min-entropy $\alpha n$, there exists such $\delta=\theta(\alpha)$ that $\pi_{1}(U)$ can be written as a convex combination of two distributions $C_{3}$ and $R$ (that is, $\pi(U)=\gamma_{3}C_{3}+\gamma R$ for some $\gamma_{3},\gamma>0,\gamma_{3}+\gamma=1$) with the following properties:
$\gamma\leq 2^{-\delta\alpha n}$ (2)
@@
$P_{v\leftarrow C_{3}}(v\in\{0,1\}^{n}\times\{0,1\}^{n}\times\{x\})\leq 2^{-(1+\delta)\alpha n}.$ (3)
@@
This of course means that $\textrm{CondDeg}_{\alpha}\leq 3-\delta+\frac{1}{\alpha n}$, which is strictly smaller than $3$ for sufficiently large $n$, so $\pi_{1}$ really has nontrivial conductance.
@@
Interestingly and importantly, in this work we managed to show that specifying more complex constructions is not possible at all. That is: every permutation with nontrivial conductance must have a character exactly like $\pi_{3}$ – so it must be a ”mix” of different concrete condensers on different coordinates (i.e. it must be a Somewhere-Condenser). This fact is described in detail in Theorem 2.
@@ 2.1 Proof of Lemma 2
To prove the statement, we need to consider two arbitrary $(2^{\alpha n})$-boxes $U=U_{1}\times\ldots\times U_{w},V=V_{1}\times\ldots\times V_{w}$ (i.e. $|U_{1}|=\ldots=|U_{w}|=|V_{1}|=\ldots=|V_{w}|=2^{\alpha n}$) and show that $|\pi(U)\cap V|\leq(2^{\alpha n})^{w-\epsilon_{1}}+\epsilon_{2}(2^{\alpha n})^{w}$. Equivalently, we will consider the uniform distribution on $U$ and prove that
$P_{u\leftarrow U}(\pi(u)\in V)\leq 2^{-\epsilon_{1}\alpha n}+\epsilon_{2}.$
@@
where $\gamma\leq\epsilon_{2}$ and $C_{i}]_{i}$ (the $i$-th coordinate of $C_{i}$) has min-entropy of at least $(1+\epsilon_{1})\alpha n$.
@@
$P_{v\leftarrow C_{i}}(v\in\{0,1\}^{n}\times\ldots\times\underset{i\text{th-position}}{\{x\}}\times\ldots\times\{0,1\}^{n})\leq 2^{-(1+\epsilon_{1})\alpha n}.$
@@
$\begin{split}P_{v\leftarrow C_{i}}(v\in V)&\leq\sum_{j=1}^{2^{\alpha n}}P_{v\leftarrow C_{i}}(v\in\{0,1\}^{n}\times\ldots\times\underset{i\text{th-position}}{\{x_{j}\}}\times\ldots\times\{0,1\}^{n})\leq\\ &\leq 2^{\alpha n}\cdot 2^{-(1+\epsilon_{1})\alpha n}=2^{-\epsilon_{1}\alpha n}.\end{split}$ (4)

### v2

@@ Abstract
In this paper, we present the first explicit examples of low-conductance permutations. The notion of conductance of permutations was introduced by Dodis et al. in ’’Indifferentiability of Confusion-Diffusion Networks‘‘, where the search for low-conductance permutations was first initiated and motivated. As part of our contribution, we not only provide these examples, but also offer a general characterization of the problem: we show that low-conductance permutations are equivalent to permutations possessing the information-theoretic properties of Multi-Source-Somewhere-Condensers, a specific variant of somewhere condensers.
@@ 1.1 Background
The starting point of this paper is the concept of permutation conductance, introduced by Yevgeniy Dodis, Martijn Stam, John Steinberger, and Tianren Liu in their seminal paper Indifferentiability of confusion-diffusion networks [11].
Their work focuses on the theoretical analysis of the indifferentiability security of confusion-diffusion networks, investigating how many rounds of certain block cipher and hash function constructions suffice to ensure adequate security. A central primitive in their analysis is a particular permutation, for which the parameter they introduced, known as conductance, is expected to be small. The results of this paper (e.g. Theorem 1) have a direct impact on these applications, as outlined in Section 1.1.2. For a detailed discussion of the motivation and the direct connections between permutations with suitable conductance and the security of confusion-diffusion based ciphers, we refer to [11]. In this work, however, the conductance of a permutation is primarily treated as a combinatorial concept, situated within the broader framework of the combinatorics of data structures.
@@
From a formal perspective, each permutation of the type $\pi:\{0,1\}^{nw}\to\{0,1\}^{nw}$ has a certain degree of conductance, depending additionally on a parameter $\alpha\in(0,1)$. This degree (denoted $\textrm{CondDeg}_{\alpha}(\pi)$) is always a real number between $1$ and $w$ and serves as a measure of how random $\pi$ is (the more random the permutation, the closer $\textrm{CondDeg}_{\alpha}$ is to $1$; in particular, a constant permutation has $\textrm{CondDeg}_{\alpha}$ equal to $w$). It turns out that permutations with low $\textrm{CondDeg}_{\alpha}$ (’’more random‘‘) are useful in the applications described in [11]. Fortunately, the authors of [11] showed (Theorem 3, App. B, probabilistic method) that almost every random permutation has (for parameters $\alpha<\frac{1}{2}-\frac{1}{nw}$, which is sufficient for their applications) $\textrm{CondDeg}_{\alpha}$ very close to $1$ (more precisely: it is no larger than $1+\frac{\log(3nw)}{\alpha n}$). On the other hand, no explicit constructions have been shown demonstrating how such permutations could be constructed in practice.
@@
In this paper we provide, for the first time, explicit examples of permutations with conductance degree less than $w$ (see Theorem 1). Our constructions are based on what we call Multi-Source-Somewhere-Condenser permutations (see Definition 5), a variant of classical somewhere condensers, such as those defined for example in Definition 1.11 of [25].
@@
Moreover, we establish a rather surprising converse result: every permutation with sufficiently low conductance degree must satisfy the Multi-Source-Somewhere-Condenser assumptions. In other words, this work provides an almost complete characterization of the problem, which is successful for nearly the entire range of parameters (see Section 3.2 for details).
1.1.1 Relations to the State-of-the-Art and Previous Results
In the literature on information theory and its applications in computer science, numerous constructions have been developed to simulate or enhance randomness. These include extractors [26, 32], condensers [30, 13], expanders [16, 23], list-decodable and list-recoverable codes [23, 29, 13], samplers [15], and pairwise independent hash functions [7, 27]. Many of these primitives are closely related or, in a certain sense, formally equivalent (see [18, 11, 28, 19, 12]), and they play a key role in managing randomness in computation. Within this framework, these constructions have numerous applications in cryptography, including universal hashing, key derivation functions, and randomness extractors for secure computation.
In this work, we show that the notion of a low-conductance permutation, as introduced in [11], can be viewed as essentially equivalent to condensers, and thus as another member of this family. Therefore, while combinatorially interesting on its own, it turns out to be equivalent to well-studied primitives. This equivalence is valuable as it enables transferring insights and parameter bounds from classical constructions to low-conductance permutations. At the same time, determining explicit parameters and limitations remains challenging, as problems in this family are inherently difficult. Overall, this perspective establishes a conceptual bridge between the combinatorial properties of permutations and information-theoretic tools, highlighting potential avenues for designing new cryptographic primitives with provable security guarantees.
1.1.2 Concrete Implications for Cryptography
A natural question is whether our results have direct implications for block cipher design. The detailed analysis of the connections between low-conductance permutations and block cipher security is thoroughly discussed in [11]. Here, we highlight the two main insights.
First, the bounds we obtain are non-trivial (see Theorem 1), though still far from the ideal case of random permutations (in the probabilistic sense), for which conductance is close to $1$. In contrast, the analysis in [11] relies on access to such nearly optimal bounds. As a result, formal guarantees can only be established for relatively stringent conductance thresholds. Unfortunately, due to the equivalence with condensers, constructing explicit examples that achieve such bounds remains a major challenge.
Second, it turns out that simple linear permutations (as originally hypothesized in [11]) are unfortunately insufficient for achieving low conductance. These functions inherently fall short of the required conductance levels because they lack the necessary condenser properties, emphasizing the need for more sophisticated constructions to approach the theoretical optimum.
Overall, while immediate implications for provably secure cipher design are limited, this work provides a rigorous combinatorial and information-theoretic foundation, clarifying the limits and potential strategies for constructing robust block ciphers.
1.1.3 Final Remarks
To conclude the introduction, we highlight how this problem exemplifies the non-obvious interconnections between several fields in theoretical computer science and mathematics. The problem originally arises in the theoretical analysis of symmetric ciphers – a relatively rare research area (see, e.g., [24, 22, 14, 8]). From this context, a notable combinatorial problem was extracted [11]. Its analysis draws heavily on techniques from information theory, particularly the theory of min-entropy extraction, and can thus also be seen as a contribution to that field (see, e.g., [5, 6, 31, 21, 20, 34, 25, 9, 10]).
It is important to note that leveraging extractor-theoretic results to tackle combinatorial problems is not entirely new; similar phenomena have been observed in various contexts (see, e.g., [17, 33, 2, 3, 4, 1]). Together, these observations illustrate why the problem under consideration is both compelling and surprisingly non-trivial, requiring insights from multiple domains.
@@ 1.2.1 Main Notion
Since the above notion is exceptionally essential for this work, to increase its readability we add an additional, intuitive understanding of this notion below preceded by one more important definition:
@@
In simple terms, $\textrm{CondDeg}_{\alpha}(\pi)$ determines how much a given permutation $\pi$ distorts all discrete $q$-boxes of dimension $w$ for $q=2^{\alpha n}$. More specifically, for each such $q$-box $U$ we look at its image $\pi(U)$ and search for another $q$-box $V$ such that the intersection $U\cap V$ is as large as possible. The maximum we can obtain in this way is given by $\textrm{CondDeg}_{\alpha}$ in the form of the following relation:
@@
In the ’’worst‘‘ (least random) case we will find such $U$ that $\pi(U)$ is just another $q$-box $V$ and then $|\pi(U)\cap V|=q^{w}$, which gives $\textrm{CondDeg}_{\alpha}=w$.
On the other hand, ’’more random‘‘ permutations should intuitively behave in such a way that regardless of the choice of $U$ and $V$, the cardinality of the set $\pi(U)\cap V$ will always be much smaller than $q^{w}$. At the same time, it can be observed that it is always possible to choose $V$ such that $|\pi(U)\cap V|$ is at least $q$. In particular, this means that we definitely have $\textrm{CondDeg}_{\alpha}(\pi)\xgeq 1$.
@@ 1.2.2 Main result
In this paper we present the first example of a permutation with nontrivial conductance degree:
@@
If $n>n_{0}$ is prime and $\epsilon\xleq\alpha\xleq 0.8$, then there exists an efficiently computable permutation $\pi:\{0,1\}^{wn}\to\{0,1\}^{wn}$ such that
$\textrm{CondDeg}_{\alpha}(\pi)\xleq w-\left\lfloor\frac{w}{3}\right\rfloor c,$
@@
The permutation $\pi:\{0,1\}^{wn}\to\{0,1\}^{wn}$ will be called a ($\alpha,\epsilon_{1},\epsilon_{2}$)-Multi-Source-Somewhere-Condenser if for any independent random variables $X_{1},X_{2},\ldots X_{w}\in\{0,1\}^{n}$ with min-entropy at least $\alpha n$, the distribution $\pi(X_{1},\ldots,X_{w})$ can be expressed as a convex combination of the distributions ($\sum_{1\xleq i\xleq w}\gamma_{i}C_{i}+\gamma R$), such that $\gamma\xleq\epsilon_{2}$ and $C_{i}]_{i}$ (i.e. the $i$-th coordinate of $C_{i}$) has min-entropy at least $(1+\epsilon_{1})\alpha n$.
Remark: the definition above is based on Definition 1.11 from [25]. However we consider several sources of the input and we limit ourselves to the study of permutations. Nevertheless, its intuitive meaning remains the same: we define a transformation, which from several sources with smaller min-entropy, leads to a convex combination of distributions, each of which (except for one component of size $\gamma$) has min-entropy on some coordinate by a factor $(1+\epsilon_{1})$ larger.
Armed with the above definitions, we are ready to formulate the following technical lemma, the proof of which can be found in Section 2.1:
@@
The formula in the above lemma means that if $\epsilon_{2}$ is negligible, then we will achieve $\textrm{CondDeg}_{\alpha}\approx w-\epsilon_{1}$.
So, to effectively use Lemma 2, we need a Multi-Source-Somewhere-Condenser with suitably strong parameters, which fortunately exists and may be found for example in the paper [5]. Specifically, it holds (in [5] this result and its proof are described as Lemma 3.14):
@@
If $\epsilon\xleq\alpha\xleq 0.8$ and $n$ is prime, then the permutation $\pi$ defined above is ($\alpha,\delta,2^{-\delta\alpha n}$)-Multi-Source-Somewhere-Condenser.
@@
If $\epsilon\xleq\alpha\xleq 0.8$, then there exists a permutation $\pi:\{0,1\}^{3n}\to\{0,1\}^{3n}$ such that
$\textrm{CondDeg}_{\alpha}(\pi)\xleq 3-c\text{ (for some }c>0\text{),}$
@@ 1.2.3 Converse Theorem
The proof is given in Section 2.2. It is also the most technically demanding part of the paper.
@@ 1.3 Intuition and Toy Examples
Consider the function $\pi_{1}(a,b,c)=(a,b,c+ab)$, where $a,b,c\in GF(2^{n})$ and $n$ is a prime number. This is of course a permutation, since for any value of $(x,y,z)$, its unique preimage is (x,y,zxy)(x,y,z-xy). We want to justify why such defined $\pi_{1}:\{0,1\}^{3n}\to\{0,1\}^{3n}$ has conductance degree $\textrm{CondDeg}_{\alpha}(\pi_{1})$ (for some fixed parameter $\alpha$) strictly less than $3$.
@@
To achieve our goal, we will use the fact that $\pi_{1}$ is a Multi-Source-Somewhere-Condenser (thus Lemma 3), and even a stronger fact (which is formulated in this way in [5]), i.e. that it is always the third (not some) coordinate of $\pi_{1}(a,b,c)$ that is the condenser, except for some negligible part of the domain. More specifically, thanks to this version of Lemma 3 we know that – since we are considering a uniform distribution on $U$ – and thus we have that $U_{1},U_{2}$ and $U_{3}$ are independent and have min-entropy $\alpha n$, there exists such $\delta=\Theta(\alpha)$ that $\pi_{1}(U)$ can be written as a convex combination of two distributions $C_{3}$ and $R$ (that is, $\pi(U)=\gamma_{3}C_{3}+\gamma R$ for some $\gamma_{3},\gamma\xgeq 0,\gamma_{3}+\gamma=1$) with the following properties:
$\gamma\xleq 2^{-\delta\alpha n}$ (2)
@@
$P_{v\leftarrow C_{3}}(v\in\{0,1\}^{n}\times\{0,1\}^{n}\times\{x\})\xleq 2^{-(1+\delta)\alpha n}.$ (3)
@@
This of course means that $\textrm{CondDeg}_{\alpha}\xleq 3-\delta+\frac{1}{\alpha n}$, which is strictly smaller than $3$ for sufficiently large $n$, so $\pi_{1}$ really has nontrivial conductance.
@@
Interestingly and importantly, in this work we managed to show that specifying more complex constructions is not possible at all. That is: every permutation with nontrivial conductance cannot be constructed in any essentially more complicated or fundamentally different way than $\pi_{3}$ – so it always must be a ’’mix‘‘ of different concrete condensers on different coordinates (i.e. it must be a somewhere condenser). This fact is described in detail in Theorem 2.
@@ 2.1 Proof of Lemma 2
To prove the statement, we need to consider two arbitrary $(2^{\alpha n})$-boxes $U=U_{1}\times\ldots\times U_{w},V=V_{1}\times\ldots\times V_{w}$ (i.e. $|U_{1}|=\ldots=|U_{w}|=|V_{1}|=\ldots=|V_{w}|=2^{\alpha n}$) and show that $|\pi(U)\cap V|\xleq(2^{\alpha n})^{w-\epsilon_{1}}+\epsilon_{2}(2^{\alpha n})^{w}$. Equivalently, we will consider the uniform distribution on $U$ and prove that
$P_{u\leftarrow U}(\pi(u)\in V)\xleq 2^{-\epsilon_{1}\alpha n}+\epsilon_{2}.$
@@
where $\gamma\xleq\epsilon_{2}$ and $C_{i}]_{i}$ (the $i$-th coordinate of $C_{i}$) has min-entropy of at least $(1+\epsilon_{1})\alpha n$.
@@
$P_{v\leftarrow C_{i}}(v\in\{0,1\}^{n}\times\ldots\times\underset{i\text{th-position}}{\{x\}}\times\ldots\times\{0,1\}^{n})\xleq 2^{-(1+\epsilon_{1})\alpha n}.$
@@
$\begin{split}P_{v\leftarrow C_{i}}(v\in V)&\xleq\sum_{j=1}^{2^{\alpha n}}P_{v\leftarrow C_{i}}(v\in\{0,1\}^{n}\times\ldots\times\underset{i\text{th-position}}{\{x_{j}\}}\times\ldots\times\{0,1\}^{n})\xleq\\ &\xleq 2^{\alpha n}\cdot 2^{-(1+\epsilon_{1})\alpha n}=2^{-\epsilon_{1}\alpha n}.\end{split}$ (4)